# IRIF, Université Paris Cité, France

### v1

@@ 0.C.2 Correlating the Weights’ Signs
$\frac{\partial^{2}_{\Delta}f(x)}{\partial^{2}_{\Delta_{0}}f(x)}=\frac{\big{|}(\Gamma_{x}^{(i-1)}\Delta)\cdot A_{k}^{(i)}\big{|}}{\big{|}(\Gamma_{x}^{(i-1)}\Delta_{0})\cdot A_{k}^{(i)}\big{|}}\coloneqq\frac{\big{|}\alpha|}{|\beta|}$
@@
The two values above are equal if and only if $\alpha$ and $\beta$ have the same sign. This additional test allows us to remove the absolute values and correlate the sign of our result for each direction $\Delta$ we take to that of the first direction $\Delta_{0}$ taken. This is why the sign recovery only has to find one sign per neuron rather than one sign per weight.
@@ 0.C.3 Solving for the Signature
$y_{m}=\frac{(\Gamma_{x}^{(i-1)}\Delta_{m})\cdot A_{k}^{(i)}}{c^{(i)}_{k}},$
where $c^{(i)}_{k}=(\Gamma_{x}^{(i-1)}\Delta_{0})\cdot A_{k}^{(i)}$. To solve for $A_{k}^{(i)}$, we build the following system of equations:
$\begin{pmatrix}\Gamma_{x}^{(i-1)}\Delta_{1}\\ \Gamma_{x}^{(i-1)}\Delta_{2}\\ \vdots\\ \Gamma_{x}^{(i-1)}\Delta_{d_{i-1}+1}\end{pmatrix}\cdot\frac{1}{c^{(i)}_{k}}\begin{pmatrix}a^{(i)}_{k,1}\\ a^{(i)}_{k,2}\\ \vdots\\ a^{(i)}_{k,d_{i-1}+1}\end{pmatrix}=\begin{pmatrix}1\\ y_{2}\\ \vdots\\ y_{d_{i-1}+1}\\ \end{pmatrix}$
@@ Appendix 0.D Deeper Critical Points Merging with Components on the Target Layer
We note $y_{1}$ and $y_{2}$ the matrix rows extracted from $x_{1}$ and $x_{2}$ respectively. We therefore have,
@@
Let’s write $A_{x_{2}}$ the matrix $A^{\prime}_{x_{2}}\circ I^{(i)}_{x_{2}}\circ\ell^{(i)}$, where $A^{\prime}_{x_{2}}=\ell^{(i+t-1)}\circ I^{(i+t-2)}_{x_{2}}\circ\dots\circ I^{(i+1)}_{x_{2}}\circ\ell^{(i+1)}$. Thus, for $c\in\mathbb{R}$,
@@
$=(\Gamma_{x_{2}}^{(i-1)}\Delta)\cdot[(\ell^{(i)\top}\circ I^{(i)\top}_{x_{2}}\circ A^{\prime\top}_{x_{2}})y_{2}]$
@@
Let’s explain the last equality. Our assumption that $x_{2}$ causes $\eta_{k}^{(i)}$ to be the only active neuron on its layer implies that $I^{(i)}$ is diagonal with all entries zero except for a one at the $k$-th diagonal position. We also have that $y_{1}$ is the $k$-th row of the matrix $\ell^{(i)}$. Therefore, $(\ell^{(i)\top}\circ I^{(i)\top}_{x_{2}}\circ A^{\prime\top}_{x_{2}})y_{2}=c\cdot y_{1}$ , where $c$ is a constant. The partial signatures $y_{1}$ and $c\cdot y_{1}$ extracted from $x_{1}$ and $x_{2}$ can be thus merged even though $x_{2}$ is on a deeper layer.
Appendix 0.E Extraction Results Using 6,000 Critical Points
Table 5: Additional results on the proportion and count of each type of unrecovered weight in models II and III, extracted with $3,000$ or $6,000$ critical points.
@@ Appendix 0.E Extraction Results Using 6,000 Critical Points
II always-off weights 00 $84$ $73.04\%$
$3,000$ CPs unreachable inactive weights 00 $12$ $10.43\%$
unreachable active weights $\leq 0.06$ $5$ $4.35\%$
$>0.06$ $3$ $2.61\%$
query-intensive weights $\leq 0.06$ $7$ $6.09\%$
$>0.06$ $4$ $3.48\%$
[…]
$6,000$ CPs unreachable inactive weights 00 $10$ $9.35\%$
unreachable active weights $\leq 0.06$ $5$ $4.67\%$
$>0.06$ $3$ $2.80\%$
query-intensive weights $\leq 0.06$ $1$ $0.93\%$
$>0.06$ $4$ $3.74\%$
[…]
$3,000$ CPs unreachable inactive weights 00 $50$ $32.26\%$
unreachable active weights $\leq 0.06$ 00 $0\%$
$>0.06$ 00 $0\%$
query-intensive weights $\leq 0.06$ $22$ $14.19\%$
$>0.06$ $3$ $1.94\%$
[…]
$6,000$ CPs unreachable inactive weights 00 $24$ $23.30\%$
unreachable active weights $\leq 0.06$ 00 $0\%$
$>0.06$ 00 $0\%$
query-intensive weights $\leq 0.06$ $6$ $5.83\%$
$>0.06$ $1$ $0.97\%$
Table 6: Additional extraction results on models II and III. Each layer is attacked independently with $3,000$ or $6,000$ critical points.
Model Signature Extraction L2 L3 L4 L5 L6 L7 L8 Entire Model
Ours
MNIST Layer coverage $100\%$ $100\%$ $100\%$ $93.73\%$ $93.73\%$ $76.32\%$ $88.58\%$ $74.12\%$
$1784-8^{(8)}-$ 1 (II) Weights Recovered $100\%$ $100\%$ $100\%$ $90.63\%$ $93.75\%$ $90.63\%$ $95.31\%$ $99.72\%$
$3,000$ critical points Time $7$m5555s $7$m37373737s $10$m41414141s $11$m46464646s $28$m29292929s $20$m38383838s $36$m46464646s -
Queries $2^{24.21}$ $2^{24.22}$ $2^{24.24}$ $2^{24.26}$ $2^{25.37}$ $2^{24.98}$ $2^{25.63}$
[…]
$1784-8^{(8)}-$ 1 (II) Weights Recovered $100\%$ $100\%$ $100\%$ $98.44\%$ $95.31\%$ $90.63\%$ $95.31\%$ $99.81\%$
$6,000$ critical points Time $14$m10101010s $15$m57575757s $28$m25252525s $29$m1111s $38$m41414141s $38$m14141414s $56$m15151515s -
Queries $2^{25.21}$ $2^{25.22}$ $2^{25.24}$ $2^{25.35}$ $2^{25.83}$ $2^{25.79}$ $2^{26.18}$
[…]
$1784-16^{(8)}-$ 1 (III) Weights Recovered $100\%$ $100\%$ $99.61\%$ $99.22\%$ $97.27\%$ $98.83\%$ $95.31\%$ $99.83\%$
$3,000$ critical points Time $26$m13131313s $28$m57575757s $39$m33333333s $1$h39393939m $2$h13131313m $1$h48484848m $3$h56565656m -
Queries $2^{24.22}$ $2^{24.24}$ $2^{24.67}$ $2^{25.65}$ $2^{26.04}$ $2^{25.79}$ $2^{26.95}$
[…]
$1784-16^{(8)}-$ 1 (III) Weights Recovered $100\%$ $99.61\%$ $100\%$ $100\%$ $99.61\%$ $99.61\%$ $98.44\%$ $99.91\%$
$6,000$ critical points Time $55$m3333s $58$m5555s $1$h4444m $2$h8888m $2$h35353535m $3$h6666m $5$h1111m -
Queries $2^{25.22}$ $2^{25.24}$ $2^{25.27}$ $2^{25.68}$ $2^{25.88}$ $2^{26.43}$ $2^{27.23}$
Table 5 shows that as the number of critical points increases, the number of query-intensive weights decreases. This suggests that some of these weights can be activated and recovered, leading to higher model fidelity. The number of dead weights also drops with more critical points, as weight plus-plus activation and neuron activation values below $0.001$ are initially considered always-off. With more critical points, these weights and neurons can become active. However, the number of unreachable active weights remains unchanged, continuing to limit full model extraction.
Table LABEL:table:result_summary_table_6000 presents the extraction results on models II and III using 6,000 critical points, compared to those using 3,000. Doubling the number of critical points generally improves both layer fidelity and weight recovery rate in deeper layers, as more query-intensive weights can be activated and recovered. Two notable exceptions are observed. First, in layer 8 of model II, the weight recovery rate is identical for both settings, but layer fidelity improves. This is because some dead weights for 3,000 critical points became active for 6,000 critical points, and were then recovered. It does not impact the weight recovery rate but slightly improves layer fidelity. Second, in layer 3 of model III, both the weight recovery rate and layer fidelity are lower with 6,000 critical points. This is due to the randomness in critical point selection. A weight with a weight plus-plus activation value of $0.012$ is only activated among the 3,000 critical points set.

### v2

@@ 0.C.2 Correlating the Weights’ Signs
$\frac{\partial^{2}_{\Delta}f(x)}{\partial^{2}_{\Delta_{0}}f(x)}=\frac{\big|A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta)\big|}{\big|A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta_{0})\big|}\coloneqq\frac{\big|\alpha|}{|\beta|}$
@@
The two values above are equal if and only if $\alpha$ and $\beta$ have the same sign. This additional test allows us to remove the absolute values and correlate the sign of our result for each direction $\Delta$ we take to that of the first direction $\Delta_{0}$ taken. This is why the sign extraction only has to find one sign per neuron rather than one sign per weight.
@@ 0.C.3 Solving for the Signature
$y_{m}=\frac{A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta_{m})}{c^{(i)}_{k}}$,
where $c^{(i)}_{k}=A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta_{0})$. To solve for $A_{k}^{(i)}$, we build the following system of equations:
$\begin{pmatrix}\Gamma_{x}^{(i-1)}\Delta_{0}\\ \Gamma_{x}^{(i-1)}\Delta_{1}\\ \Gamma_{x}^{(i-1)}\Delta_{2}\\ \vdots\\ \Gamma_{x}^{(i-1)}\Delta_{d_{i-1}-1}\end{pmatrix}\cdot\frac{1}{c^{(i)}_{k}}\begin{pmatrix}a^{(i)}_{k,0}\\ a^{(i)}_{k,1}\\ a^{(i)}_{k,2}\\ \vdots\\ a^{(i)}_{k,d_{i-1}-1}\end{pmatrix}=\begin{pmatrix}1\\ y_{1}\\ y_{2}\\ \vdots\\ y_{d_{i-1}-1}\\ \end{pmatrix}$
@@ Appendix 0.D Deeper Critical Points Merging with Components on the Target Layer
This indicates that a strict association between deeper critical points and incorrect components can lead to mislabelling. We note $y_{1}$ and $y_{2}$ the matrix rows extracted from $x_{1}$ and $x_{2}$ respectively. We therefore have,
@@
Let’s write $A_{x_{2}}$ the matrix $A^{\prime}_{x_{2}}\circ I^{(i)}_{x_{2}}\circ A^{(i)}$, where $A^{\prime}_{x_{2}}=A^{(i+t-1)}\circ I^{(i+t-2)}_{x_{2}}\circ\dots\circ I^{(i+1)}_{x_{2}}\circ A^{(i+1)}$. Thus, for $c\in\mathbb{R}$,
@@
$=(\Gamma_{x_{2}}^{(i-1)}\Delta)\cdot[(A^{(i)\top}\circ I^{(i)\top}_{x_{2}}\circ A^{\prime\top}_{x_{2}})y_{2}]$
@@
Let’s explain the last equality. Our assumption that $x_{2}$ causes $\eta_{k}^{(i)}$ to be the only active neuron on its layer implies that $I^{(i)}$ is diagonal with all entries zero except for a one at the $k$-th diagonal position. We also have that $y_{1}$ is the $k$-th row of the matrix $A^{(i)}$. Therefore, $(A^{(i)\top}\circ I^{(i)\top}_{x_{2}}\circ A^{\prime\top}_{x_{2}})y_{2}=c\cdot y_{1}$ , where $c$ is a constant. The partial signatures $y_{1}$ and $c\cdot y_{1}$ extracted from $x_{1}$ and $x_{2}$ can be thus merged even though $x_{2}$ is on a deeper layer.
On a side note, when this happens, the sign of all neurons on layer $i$ can be recovered immediately with no queries or computations as we have an input $x^{*}$ for which all neurons on the layer must be inactive (first order derivatives in $\mathcal{P}_{x^{*}}$ are equal to zero). We must have $\eta^{(i)}_{k}\circ F^{(i-1)}(x^{*})<0$, for all neurons on the layer, so we choose the sign of each $\eta^{(i)}_{k}$ to match this condition.
Appendix 0.E Noise for Large Networks
In this section, we show that recovered partial neurons from deeper layers correspond to very different mathematical expressions from neurons on the target layer, leading to the possibility of identifying deeper critical point at no cost using statistical arguments. Indeed, we can view each recovered weight as a random variable.
Let’s start with a neuron $\eta^{(i+1)}_{j}$ on the $i+1$’th layer. The equation of its hyperplane is given by:
$A^{(i+1)}_{j}\cdot F^{(i)}(x)+b^{(i+1)}_{j}=0$
Which is, rewritten according to layer i1i-1,
$A^{(i+1)}_{j}\cdot\left(I^{(i)}A^{(i)}F^{(i-1)}(x)+I^{(i)}b^{(i)}\right)+b^{(i+1)}_{j}=0$
$\sum_{k=1}^{d_{i+1}}a^{(i+1)}_{j,k}\cdot\left[I^{(i)}A^{(i)}F^{(i-1)}(x)\right]_{k}+\mathcal{B}=0$
$\sum_{k=1}^{d_{i+1}}a^{(i+1)}_{j,k}\cdot\left(\sum_{l=1}^{d_{i}}I^{(i)}_{k,k}\cdot a^{(i)}_{k,l}\cdot[F^{(i-1)}(x)]_{l}\right)+\mathcal{B}=0$
Expressing directly in terms of outputs of the extracted layer i1i-1:
$\sum_{l=1}^{d_{i}}\left(\sum_{k=1}^{d_{i+1}}a^{(i+1)}_{j,k}\cdot I^{(i)}_{k,k}\cdot a^{(i)}_{k,l}\right)[F^{(i-1)}(x)]_{l}+\mathcal{B}=0$
and so, if it is not made inactive by a ReLU on layer i1i-1, the first weight extracted is:
$\sum_{k=1}^{d_{i+1}}a^{(i+1)}_{j,k}\cdot I^{(i)}_{k,k}\cdot a^{(i)}_{k,1}$
And therefore by a quick induction, the first extracted weight of a neuron $\eta_{j}^{(n)}$ when recovering layer $i$ correspond to all the active paths between the weight we think we are extracting and $\eta^{(n)}_{j}$:
$\sum_{k_{n}=1}^{d_{n}}a^{(n)}_{j,k_{n}}\cdot I^{(n-1)}_{k_{n},k_{n}}\cdot(\sum_{k_{n-1}=1}^{d_{n-1}}a^{(n-1)}_{k_{n},k_{n-1}}\cdot I^{(n-2)}_{k_{n-1},k_{n-1}}$
$\cdot(\dots(\sum_{k_{i+1}=1}^{d_{i+1}}a^{(i+1)}_{k_{i+2},k_{i+1}}\cdot I^{(i)}_{k_{i+1},k_{i+1}}\cdot a^{(i)}_{k_{i+1},1})\dots))$
As an example, suppose for simplicity that every layer has width $d$, that for any given critical point we expect half of the neurons to be active, and that all the weights of the network have the same mean and variance. Let’s find the distribution of our extracted $w^{(n)}$ when attempting to find layer $i$. Let’s denote the set of paths as $P=\{(k_{i+1},k_{i+2},\dots,k_{n},j)|\forall m,k_{m}\in\{1,\dots,d\}\}$. Then,
$w^{(n)}=\sum_{p\in P}a^{(i)}_{k_{i}+1,1}\cdot\prod_{l=i}^{n-1}(a^{(l+1)}_{p_{l-i+2},p_{l-i+1}}\cdot I^{(l)}_{p_{l-i+1},p_{l-i+1}}),\text{where }p_{i}\text{ is the }i\text{th entry of }p$
Let’s write $L$ for nin-i, and $\mathcal{C}_{L}$ for $\frac{d^{L}}{2^{L}}$ then we have that,
$\mathbb{E}[w^{(n)}]$ $=d^{L}\cdot\mathbb{E}[w^{(i)}]\cdot(\mathbb{E}[w^{(i)}]\cdot\frac{1}{2})^{L}$
$=\mathcal{C}_{L}\cdot\mathbb{E}[w^{(i)}]^{L+1}$
and,
$\mathbb{E}[(w^{(n)})^{2}]$ $=\mathbb{E}[\sum_{p\in P}\sum_{p^{\prime}\in P}(a^{(i)}_{k_{i}+1,1})^{2}\cdot\prod_{l=i}^{n-1}a^{(l+1)}_{p_{l-i+2},p_{l-i+1}}$
$\cdot I^{(l)}_{p_{l-i+1},p_{l-i+1}}\cdot a^{(l+1)}_{p^{\prime}_{l-i+2},p^{\prime}_{l-i+1}}\cdot I^{(l)}_{p^{\prime}_{l-i+1},p^{\prime}_{l-i+1}}$]
Let’s separate between $p=p^{\prime}$ and $p\neq p^{\prime}$. For $p=p^{\prime}$, we have just as above, $d^{L}$ paths, $I^{(l)}_{p^{\prime},p^{\prime}}=I^{(l)}_{p,p}$ and $a^{(l)}_{p^{\prime}}=a^{(l)}_{p}$. For $p\neq p^{\prime}$, we have $d^{2L}-d^{L}$ possible path combinations, $I^{(l)}_{p^{\prime},p^{\prime}}\neq I^{(l)}_{p,p}$ and $a^{(l)}_{p^{\prime}}\neq a^{(l)}_{p}$. By assumption $\mathbb{E}[I^{(l)}]=\frac{1}{2}$. We therefore get summing the $p=p^{\prime}$ paths with the $p\neq p^{\prime}$ paths,
$\mathbb{E}[(w^{(n)})^{2}]=d^{L}\cdot\mathbb{E}[(w^{(i)})^{2}]\cdot(\mathbb{E}[(w^{(i)})^{2}]\cdot\frac{1}{2})^{L}+(d^{2L}-d^{L})\cdot\mathbb{E}[w^{(i)}]^{2}\cdot(\mathbb{E}[w^{(i)}]^{2}\cdot\frac{1}{2^{2}})^{L}$
Let’s simplify notation and set $\mathbb{E}[w^{(i)}]=\mu$, $\mathbb{V}(w^{(i)})=\sigma^{2}$ for all $i$ using our assumption,
$\mathbb{E}[(w^{(n)})^{2}]$ $=d^{L}\cdot(\sigma^{2}+\mu^{2})\cdot(\frac{\sigma^{2}+\mu^{2}}{2})^{L}+(d^{2L}-d^{L})\cdot\mu^{2}\cdot(\frac{\mu}{2})^{2L}$
$=\mathcal{C}_{L}\left[(\mathcal{C}_{L}-1)\mu^{2L+2}+(\sigma^{2}+\mu^{2})^{L+1}\right]$
and therefore,
$\mathbb{V}(w^{(n)})$ $=\mathbb{E}[(w^{(n)})^{2}]-\mathbb{E}[w^{(n)}]^{2}$
$=\mathcal{C}_{L}\left[(\mathcal{C}_{L}-1)\mu^{2L+2}+(\sigma^{2}+\mu^{2})^{L+1}\right]-\mathcal{C}_{L}^{2}\mu^{2L+2}$
$=\mathcal{C}_{L}[(\sigma^{2}+\mu^{2})^{L+1}-\mu^{2(L+1)}]$
Therefore for a given statistical distribution of $w^{(i)}$, the variance and the mean of weights recovered from a neuron on a deeper layer depends on how much deeper the neuron is compared to the target layer (L=niL=n-i), as well as the dimension of each layer. In the attack, a partial neuron is extracted up to a scalar. Our experiments show that despite this scalar and underdetermined systems, variations in variance are enough to identify a large proportion of points on deeper layers, see Fig. 13. Being able to identify points on deeper layers leads to the speedup discussed in Section 3.2, see Table 6.
L4 points kept (%) 100 99 95 90 80 70 60 50 20 10
L4 points proportion (%) 11.8 14.5 16.6 17.6 19.2 20.2 21.1 22.2 26.5 24.5
Expected min speedup 1 1.54 2.00 2.27 2.69 2.96 3.26 3.58 5.10 4.37
@@ Appendix 0.E Noise for Large Networks
Figure 13: For each layer, the number of critical points we discard on a $784\!-\!256^{(16)}\!-\!1$ network attacked with 100,000 points by imposing an upper bound on variance that would keep 99%, 90% and 50% of points on layer 4, respectively in purple, green and red. As implied by the mathematical analysis, critical points from deeper layers are easier to identify: the percentage of points on deeper layers kept is always lower than that of layer 4.
Appendix 0.F Impact of Scaling Factors on Precision
In this section, we give an example showing how large disparities among scaling factors degrade the precision of signature recovery. Without normalization, the scaling factors for layer 2 in a $784-8^{(8)}-1$ MNIST model are
(161.71,32.11,23.14,1.07,0.81,36.64,0.56,0)(161.71,32.11,23.14,1.07,0.81,36.64,0.56,0)
where values are rounded to two decimal places for simplicity. The ratio between the largest factor (161.71) and the smallest factor (0.56) is 288.77.
Next, we can identify a critical point for the 8th neuron in layer 3, for which the recovered weights are imprecise. The true weights for this neuron are
(0.51,0.47,0.55,0.37,0.28,0.63,0.23,0.39)(0.51,0.47,-0.55,0.37,-0.28,0.63,-0.23,-0.39)
with relative ratios (normalized to the first element)
(1,0.92,1.08,0.73,0.56,1.25,0.46,0.77)(1,0.92,-1.08,0.73,-0.56,1.25,-0.46,-0.77)
Without normalization, the recovered weights are
$(-6.33\times 10^{-4},-3.04\times 10^{-3},4.99\times 10^{-3},-7.29\times 10^{-2},6.85\times 10^{-2},-3.46\times 10^{-3},0,0)$
After multiplication by the scaling factors, they become
$(-1.02\times 10^{-1},-9.77\times 10^{-2},1.16\times 10^{-1},-7.79\times 10^{-2},5.56\times 10^{-2},-1.27\times 10^{-1},0,0)$
with relative ratios (normalized to the first element)
(1,0.95,1.13,0.76,0.54,1.24,0,0)(1,0.95,-1.13,0.76,-0.54,1.24,0,0)
After applying normalization, the recovered weights are
$(-6.34\times 10^{-4},-2.95\times 10^{-3},4.78\times 10^{-3},-7.04\times 10^{-2},7.12\times 10^{-2},-3.51\times 10^{-3},0,0)$
After multiplication by the scaling factors, they become
$(-1.03\times 10^{-1},-9.46\times 10^{-2},1.11\times 10^{-1},-7.52\times 10^{-2},5.78\times 10^{-2},-1.28\times 10^{-1},0,0)$
with relative ratios (normalized to the first element)
(1,0.92,1.08,0.73,0.56,1.25,0,0)(1,0.92,-1.08,0.73,-0.56,1.25,0,0)
With normalization, the recovered weights exhibit perfect relative ratios (0.F), matching the true ratios (0.F). In contrast, without normalization, the recovered weights are imprecise, as reflected in the relative ratios (0.F).
Appendix 0.G Increasing the Rank of SOE
In this section, we analyze the rank of our improved SOE system, which stacks the SOE systems of multiple points. Consider $n$ points $\{x_{j}\}_{j\in[1,n]}$ for sign extraction on layer $i$. These points share the same activation pattern on later layers, denoted as $G_{x_{j}}^{(i+1)}$ (simplified as $G^{(i+1)}$), but trigger many different activation patterns for $F_{x_{j}}^{(i-1)}$. For each point $x_{j}$, we can construct a SOE system:
$\{G^{(i+1)}I^{(i)}A^{(i)}F_{x_{j}}^{(i-1)}\Delta_{k_{j}}=f(x_{j}+\Delta_{k_{j}})-f(x_{j})\}_{k_{j}}$
by taking multiple input directions $\Delta_{k_{j}}$, where our unknown is $G^{(i+1)}I^{(i)}$. Stacking these systems yields a larger SOE system:
$G^{(i+1)}I^{(i)}\begin{pmatrix}\{A^{(i)}F_{x_{1}}^{(i-1)}\Delta_{k_{1}}\}_{k_{1}}\\ \{A^{(i)}F_{x_{2}}^{(i-1)}\Delta_{k_{2}}\}_{k_{2}}\\ \vdots\\ \{A^{(i)}F_{x_{n}}^{(i-1)}\Delta_{k_{n}}\}_{k_{n}}\end{pmatrix}=\begin{pmatrix}B_{1}\\ B_{2}\\ \vdots\\ B_{n}\end{pmatrix}=\begin{pmatrix}\{f(x_{1}+\Delta_{k_{1}})-f(x_{1})\}_{k_{1}}\\ \{f(x_{2}+\Delta_{k_{2}})-f(x_{2})\}_{k_{2}}\\ \vdots\\ \{f(x_{n}+\Delta_{k_{n}})-f(x_{n})\}_{k_{n}}\\ \end{pmatrix}$
where each block matrix $\{A^{(i)}F_{x_{j}}^{(i-1)}\Delta_{k_{j}}\}_{k_{j}}$ is denoted as $B_{j}$.
In this stacked SOE system, our focus is on the rank of
$C=\begin{pmatrix}B_{1}\\ B_{2}\\ \vdots\\ B_{n}\end{pmatrix}$
as rank deficiency determines how many inactive neurons with high confidence must be recovered using neuron wiggle.
First, we give a trivial bound for $\text{rank}(C)$. Since $C$ contains all rows of $\{B_{j}\}_{j\in[1,n]}$, $\text{rank}(C)$ is at least the largest rank among the individual blocks: $\max_{1\leq j\leq n}\text{rank}(B_{j})$. On the other hand, the row space of $C$ is spanned by the row spaces of all $\{B_{j}\}_{j\in[1,n]}$, so $\text{rank}(C)$ cannot exceed the sum of the ranks of these blocks: $\Sigma_{j=1}^{n}\text{rank}(B_{j})$. Therefore, a loose bound for $\text{rank}(C)$ is:
$\max_{1\leq j\leq n}\text{rank}(B_{j})\leq\text{rank}(C)\leq\Sigma_{j=1}^{n}\text{rank}(B_{j})$
Next, we present a method for exact rank calculation. Since the row space of $C$ is spanned by the row spaces of all $\{B_{j}\}_{j\in[1,n]}$, its rank is the sum of the ranks of the individual blocks minus the dimension of any shared row spaces, to avoid double-counting. This yields
$\text{rank}(C)=$ $\Sigma_{j=1}^{n}\text{rank}(B_{j})-\Sigma_{1\leq j<k\leq n}\dim(R(B_{j})\cap R(B_{k}))$
$+\Sigma_{1\leq j<k<l\leq n}\dim(R(B_{j})\cap R(B_{k})\cap R(B_{l}))$
$-\cdots+(-1)^{n-1}\dim(R(B_{1})\cap R(B_{2})\cap\cdots\cap R(B_{n}))$
where $\dim(\cdot)$ denotes the dimension of a space, and $R(B_{j})$ represents the row space spanned by $B_{j}$.